OpenStudy (jusaquikie):

how many chloride ions arein 25 mL of an aquious solution of 0.224 M MgCl2

OpenStudy (aaronq):

find how many moles you have, then knowing that 1 mole = 6.023x10^23 atoms/ions/molecules, find how many ions you got

OpenStudy (jusaquikie):

i know the MM of MgCl2= 95.21 and i know that there are 6.022 ions per mole but i missing the inbetween i think. i converted it to grams but don't think i'm going the right direction.

OpenStudy (jusaquikie):

so to get the moles of Cl i would take the MM of Cl2 divided by the MM of MgCL2?

OpenStudy (aaronq):

you're working with molarity (M), so use the formula for molarity and plug stuff n solve for moles

OpenStudy (aaronq):

no, you don't need the molar mass for this problem, but you do need to know that MgCl2 dissociates into Mg^2+ + 2Cl^-

OpenStudy (aaronq):

so the amount of ions will be 3 times that of the moles

OpenStudy (aaronq):

you have molarity=moles/L of solvent M=n/V 0.224 M = n/0.025 L solve for n, then multiply times avogadros number, and you know that is dissociates into 3 ions, so multiply by 3

OpenStudy (jusaquikie):

Thank you this makes sense now.

OpenStudy (aaronq):

no prob dude